# CITER: Collaborative Inference for Efficient Large Language Model Decoding with Token-Level Routing

### v4

@@ Abstract
Large language models have achieved remarkable success in various tasks but suffer from high computational costs during inference, limiting their deployment in resource-constrained applications. To address this issue, we propose a novel Collaborative Inference with Token-lEvel Routing (CITER) framework that enables efficient collaboration between small and large language models (SLMs & LLMs) through a token-level routing strategy. Specifically, CITER routes non-critical tokens to an SLM for efficiency and routes critical tokens to an LLM for generalization quality. We formulate router training as a policy optimization, where the router receives rewards based on both the quality of predictions and the inference costs of generation. This allows the router to learn to predict token-level routing scores and make routing decisions based on both the current token and the future impact of its decisions. To further accelerate the reward evaluation process, we introduce a shortcut which significantly reduces the costs of the reward estimation and improving the practicality of our approach. Extensive experiments on five benchmark datasets demonstrate that CITER reduces the inference costs while preserving high-quality generation, offering a promising solution for real-time and resource-constrained applications. Our data and code are available at https://github.com/aiming-lab/CITER.
@@ 1 Introduction
Large language models (LLMs) have revolutionized various natural language processing tasks, from machine translation to context summarization and question answering (Coleman et al., 2024; Kamalloo et al., 2024; Eniser et al., 2024; He et al., 2024). However, their impressive performance comes with a substantial computational costs, particularly during inference. As these models grow in size, the costs of inference becomes a significant barrier to their practical deployment, especially in real-time applications. Therefore, there is an increasing need to reduce inference costs without compromising the quality of the generated outputs.
@@
To address this challenge, we present a novel framework, namely Collaborative Inference with Token-lEvel Routing (CITER). CITER introduces a token-level router which determines either LLM or SLM is used to generate each token. Specifically, many tokens in the response that are not important to the final prediction, can be routed and generated by SLM to reduce inference costs, while the LLM can be reserved to generate important tokens only. We propose optimizing this router using pairwise data by reinforcement learning. The objective of the optimization is to minimize the inference costs while preserving output quality. By employing this formulation, the router learns to predict token-level routing scores and make routing decisions not only based on the current token but also considering the impact of these decisions on future tokens. In order to further accelerate the estimation of the reward function defined by the accuracy of the response, we present a surrogate reward function as a shortcut, where the predictions from the SLM and LLM are leveraged to estimate the final reward without completing the whole generation process, which significantly accelerates the training process. Through this framework, the router can optimize its token-level routing policy by RL, enabling the collaboration of SLM and LLM for effective and efficient autoregressive generation.
Our primary contribution is CITER, which reduces inference costs by employing a token-level router to select the appropriate model to generate each token. Experiments on five benchmark datasets demonstrate the effectiveness of our approach, achieving a reduction in LLM inference costs while maintaining high output accuracy, with comparable accuracy while making up to 30% fewer inference costs or delivering up to a 25% improvement in accuracy with the same cost compared to Co-LLM (Shen et al., 2024). Furthermore, our experiments in the ablation study also demonstrate that token-level routing offers more flexibility to achieve more promising results compared to query-level routing and that considering the long-term impact of routing decisions significantly boosts acceleration performance.
Notations. We denote $\pi_{\bm{\theta}}$ as the policy model parameterized by $\bm{\theta}$, $x_{i}$ as the $i$-th token in the input prompt, and $y_{j}$ as the $j$-th token of the output response. We denote $\operatorname{\mathds{1}}[\cdot]$ as the indicator function and $\oplus$ as the concatenate operation. All other notations are defined prior to their first usage.
@@ 2 Collaborative Inference with Token-lEvel Routing (CITER)
In this section, we describe our Collaborative Inference with Token-lEvel Routing (CITER) framework that uses token-level routing to accelerate the inference of language models. As illustrated in Figure 1, in CITER, we introduce a router to facilitate collaborative inference between a powerful but computationally expensive LLM and a fast but potentially inaccurate SLM. Specifically, the router is used to predict the token-level routing score for each token, and a predefined threshold $\tau$ is used to determine which model should generate this token. The key challenge of our framework is the router training process. To feed the router with the knowledge on making the global optimal routing decisions not only based on the accuracy of the current token but also the long-term impact of its decision, we formulate the router training process as a preference-based policy optimization task, aiming to minimize the inference costs while maintaining the generation quality. To be more specific, we first formulate the RL problem and derive the reward function as token-wise routing preference, which should be computed to collect during the router training process. Subsequently, we introduce a shortcut for the reward function estimation, leveraging both the SLM and LLM’s prediction to estimate the reward, to significantly accelerate the collection process of the token-wise routing preference. Finally, we propose an iterative router training process to mitigate the potential inconsistencies of routing decisions in the preference collection phase and deployment. In the rest of this section, we will outline router training and collaborative inference processes in detail.
@@ 2.1 Reinforcement Learning for Router Optimization
where $\mu$ is the prior policy intuitively related to the cost difference for evaluating LLM or SLM. The expectation is taken over the randomness of language models, policy $\pi$ and the prompt $\mathbf{s}_{0}$.
@@ 2.2 Preference-based Token-level Policy Optimization
where $\sigma(z):=1/(1+\exp(-z))$ is the sigmoid function. Following Rafailov et al. (2024), we have
$Q_{h}^{*}(\mathbf{s}_{h},a_{L})-Q_{h}^{*}(\mathbf{s}_{h},a_{S})=$ $\beta\log\frac{\pi_{h}(a_{L}|\mathbf{s}_{h})}{\mu(a_{L})}-\beta\log\frac{\pi_{h}(a_{S}|\mathbf{s}_{h})}{\mu(a_{S})}$
$=$ $\beta\log\frac{\pi_{h}(a_{L}|\mathbf{s}_{h})}{\pi_{h}(a_{S}|\mathbf{s}_{h})}-\beta\log\frac{\mu(a_{L})}{\mu(a_{S})},$ (4)
@@
$\Pr(a_{S}\succ a_{L}|\mathbf{s}_{h})$ $=\frac{1}{1+\frac{\pi_{h}(a_{L}|\mathbf{s}_{h})}{\pi_{h}(a_{S}|\mathbf{s}_{h})}}=\pi_{h}(a_{S}|\mathbf{s}_{h}),$
where the latter equation is due to the fact that $\pi_{h}(a_{S}|\mathbf{s}_{h})+\pi_{h}(a_{L}|\mathbf{s}_{h})=1$. Therefore, given a sequence of token $\mathbf{s}_{h}$, once we have labeled the preference $\operatorname{\mathds{1}}[a_{S}\succ a_{L}|\mathbf{s}_{h}]$, $\pi_{h}(a_{S}|\mathbf{s}_{h})$, the routing agent $\pi$ can be learned by minimizing the cross-entropy loss
@@
where $\operatorname{\mathds{1}}[a_{L}\succ a_{S}|\mathbf{s}_{h}]$ indicates the event where using large language model is preferred at state $\mathbf{s}_{h}$.
@@
8:           if $M_{S}(\mathbf{s}_{h})=y^{*}_{h+1}$ then Set $p_{h}=1$ /* Case 1. $a_{S}\succ a_{L}$/
9:           if $M_{L}(\mathbf{s}_{h})=y^{*}_{h+1}$ then Set $p_{h}=0$ /* Case 2. $a_{L}\succ a_{S}$/
@@
2.3 Acquiring Token-level Routing Preference
In this subsection, we describe our strategy to determine the preference label $\operatorname{\mathds{1}}[a_{L}\succ a_{S}|\mathbf{s}_{h}]$. For a state $\mathbf{s}_{h}$, we first generate the next token $y_{h+1}$ with the small language model and then complete the whole trajectory $\mathbf{s}_{H}$ until the generation ends with <EOS> using the routing policy $\pi$. Compared to equation 1, the reward collected on this trajectory $\mathbf{s}_{H}$ is an unbiased estimation of $Q_{h}^{\pi}(\mathbf{s}_{h},a_{S})$. Intuitively, if using the small language model in the current step $h$ can generate the correct response $\mathbf{s}_{H}$, then the small language model is preferred. Otherwise, we assign $a_{L}\succ a_{S}$ and assume that the large language model can generate the correct answer, as implemented in Line 10 in Algorithm 1.
However, generating and evaluating the final response $\mathbf{s}_{H}$ might be expensive or even infeasible when $H$ is large. In order to further accelerate the token-level routing preference label, we introduce a shortcut by leveraging the ground truth response $\mathbf{s}^{*}_{H}$ provided in the dataset. As Line 8 in Algorithm 1 suggests, if the next token $y_{h+1}^{S}$ generated by the small language model is exactly the same as the ground-truth token $y^{*}_{h+1}$, we assign $a_{S}\succ a_{L}$ since the behavior of the small language model is good enough to match the ground-truth model. In the case where the next token generated by the small language model does not match the ground truth, as carried out in Line 9, we check the next token generated by the large language model $y_{h+1}^{L}$ and assign $a_{L}\succ a_{S}$ if $y_{h+1}^{L}=y_{h+1}^{*}$. Otherwise we will go back to the aforementioned case to evaluate the completed generated trajectory as conducted in Line 10 in Algorithm 1. We would like to highlight that only when both models fail to generate the correct token $y^{*}_{h}$ based on ground truth context, the full response generation is required to compute the reward. This shortcut allows us to obtain routing preferences for most tokens without generating the full response. Empirically, we find that about $80\%\sim 90\%$ of the tokens can be correctly predicted by either the SLM or LLM, which makes the shortcut significantly reduce the inference costs of estimating the reward function.
2.4 Proposed Algorithm
The inference algorithm CITER is presented in Algorithm 2. In detail, CITER uses a deterministic policy where it chooses the small language model when $\pi(a_{S}|\mathbf{s})\geq\pi(\allowbreak|\mathbf{s})$ (i.e., $\pi(a_{S}|\mathbf{s})\geq\frac{1}{2}$) and vice versa. To further investigate the balance between efficiency and precision by collaborating with LLM and SLM, we introduce another layer of prior policy $(\rho(a_{S}),\rho(a_{L}))$. Thus, the deterministic rule of selecting the SLM from the posterior policy distribution $\pi^{\prime}(a|\mathbf{s})\propto\pi(a|\mathbf{s})\rho(a)$ is that
Remark 2.1.
We maintain separate KV caches for SLM and LLM. When the CITER algorithm switches between them, the KV cache from the previous model is preserved, allowing it to be reused when switching back. This eliminates the need for redundant computations, improving efficiency.
@@ 2.4 Proposed Algorithm
5:     Generate next token and let $\tilde{\mathbf{s}}=\mathbf{s}\oplus\{M(\tilde{\mathbf{s}})\}$
@@ 3 Experiments
In this section, we evaluate the performance of CITER aiming to answer the following questions: (1) Compared with the previous works on speeding up the inference of LLM, how does our framework perform in terms of the inference costs and the quality of the generated response? (2) Does the components we proposed in our framework boost the performance of the router? (3) Does the iterative training process of the router improve the performance of our framework? (4) How does the performance of our framework change with the size of the LLM? (5) Can the router distinguish the critical and non-critical tokens correctly?
@@ 3.1 Experimental Setup
Table 1: The statistics of our evaluation datasets. The commonsense QA dataset and MMLU-Professional Psychology dataset are denoted as CS QA and MMLU-PP, respectively
@@
MMLU-PP Psychology Multi-choice 4 612 69
@@
Evaluation. To evaluate the performance of CITER and the baseline methods, we use the accuracy of responses to reflect response quality and define the inference cost as the average amount of data transformation per sample since LLM generation is primarily memory-bound. Details on the data transformation calculations and an illustration of the memory-bound nature of LLM generation are provided in Appendix E. Additionally, for both query-level routing methods and token-level routing methods, a threshold $\tau$ is applied in each method to balance the trade-off between leveraging the LLM to improve the response quality or offloading to the SLM to reduce the overall inference costs. We then plot the accuracy curve versus the average amount of data transformation per sample to illustrate the acceleration performance of both CITER and the baselines. The optimal point is located in the top-left corner of the curve, corresponding to the highest accuracy with the lowest costs.
@@
Implementation Details We implement our framework using the Hugging Face Transformers library (Wolf et al., 2020). For the SLM and LLM, we utilize Qwen2-1.5b and Qwen2-72b, respectively. The router is implemented as a multilayer perceptron (MLP) network with three hidden layers, ReLU activation (Agarap, 2019), BatchNorm normalization (Ioffe & Szegedy, 2015), and a $0.1$ dropout rate. It is trained using the Adam optimizer (Kingma & Ba, 2017) with a learning rate of $1\times 10^{-7}$, betas of $(0.9,0.99)$, and no weight decay. Training is performed on a single NVIDIA H100 GPU with a batch size of $80$. The iterative training process runs for $2$ rounds.

### v6

@@ Abstract
Large language models have achieved remarkable success in various tasks but suffer from high computational costs during inference, limiting their deployment in resource-constrained applications. To address this issue, we propose a novel Collaborative Inference with Token-lEvel Routing (CITER) framework that enables efficient collaboration between small and large language models (SLMs & LLMs) through a token-level routing strategy. Specifically, CITER routes non-critical tokens to an SLM for efficiency and routes critical tokens to an LLM for generalization quality. We formulate router training as a policy optimization, where the router receives rewards based on both the quality of predictions and the inference costs of generation. This allows the router to learn to predict token-level routing scores and make routing decisions based on both the current token and the future impact of its decisions. To further accelerate the reward evaluation process, we introduce a shortcut which significantly reduces the costs of the reward estimation and improving the practicality of our approach. Extensive experiments on five benchmark datasets demonstrate that CITER reduces the inference costs while preserving high-quality generation, offering a promising solution for real-time and resource-constrained applications.
@@ 1 Introduction
Large language models (LLMs) have revolutionized various natural language processing tasks, from machine translation to context summarization and question answering (Coleman et al., 2024; Kamalloo et al., 2024; Eniser et al., 2024). However, their impressive performance comes with a substantial computational costs, particularly during inference. As these models grow in size, the costs of inference becomes a significant barrier to their practical deployment, especially in real-time applications. Therefore, there is an increasing need to reduce inference costs without compromising the quality of the generated outputs.
@@
To address this challenge, we present a novel framework, namely Collaborative Inference with Token-lEvel Routing (CITER). CITER introduces a token-level router which determines either LLM or SLM is used to generate each token. Specifically, many tokens in the response that are not important to the final prediction, can be routed and generated by SLM to reduce inference costs, while the LLM can be reserved to generate important tokens only. We propose optimizing this router using pairwise data by reinforcement learning, with the objective of minimizing the inference costs while preserving output quality. By employing this formulation, the router learns to predict token-level routing scores and make routing decisions not only based on the current token but also considering the impact of these decisions on future tokens. In order to further accelerate the estimation of the reward function defined by the accuracy of the response, we present a surrogate reward function as a shortcut, where the predictions from the SLM and LLM are leveraged to estimate the final reward without completing the whole generation process, accelerating the training process significantly. Through this framework, we enable the collaboration of SLM and LLM for effective and efficient autoregressive generation.
Our primary contribution is CITER, which reduces inference costs by employing a token-level router to select the appropriate model to generate each token. Experiments on five benchmark datasets demonstrate the effectiveness of our approach, making up to 30% fewer inference costs with comparable accuracy or delivering up to a 25% improvement in accuracy with the same cost compared to Co-LLM (Shen et al., 2024). Furthermore, our experiments in the ablation study also demonstrate that token-level routing offers more flexibility to achieve more promising results compared to query-level routing and that considering the long-term impact of routing decisions significantly boosts performance.
Notations. We denote $\pi_{\bm{\theta}}$ as the policy model parameterized by $\bm{\theta}$, $x_{i}$ as the $i$-th token in the input prompt, $y_{j}$ as the $j$-th token of the output response, $\operatorname{\mathds{1}}[\cdot]$ as the indicator function and $\oplus$ as the concatenate operation. All other notations are defined prior to their first usage.
@@ 2 Collaborative Inference with Token-lEvel Routing (CITER)
In this section, we describe our Collaborative Inference with Token-lEvel Routing (CITER) framework that uses token-level routing to accelerate the inference of language models. As illustrated in Figure 1, we introduce a router to facilitate collaborative inference between a powerful but computationally expensive LLM and a fast but potentially inaccurate SLM. Specifically, the router is used to predict the token-level routing score for each token, and a predefined threshold $\tau$ is used to determine which model should generate this token. The key challenge of our framework is the router training process. To feed the router with the knowledge on making the global optimal routing decisions not only based on the accuracy of the current token but also the long-term impact of its decision, we formulate the router training process as a preference-based policy optimization task, aiming to minimize the inference costs while maintaining the generation quality. To be more specific, we first formulate the RL problem and derive the reward function as token-wise routing preference, which should be computed to collect during the router training process. Subsequently, we introduce a shortcut for the reward function estimation, leveraging both the SLM and LLM’s prediction to estimate the reward, significantly accelerating the collection process of the token-wise routing preference. Finally, we propose an iterative router training process to mitigate the potential inconsistencies of routing decisions in the preference collection phase and deployment. In the rest of this section, we will outline router training and collaborative inference processes in detail.
@@ 2.1 Reinforcement Learning for Router Optimization
where $\mu$ is the prior policy related to the cost difference for evaluating LLM or SLM. The expectation is taken over the randomness of language models, policy $\pi$ and the prompt $\mathbf{s}_{0}$.
@@ 2.2 Preference-based Token-level Policy Optimization
where $\sigma(z)$ is the sigmoid function. Following Rafailov et al. (2024), we have
$Q_{h}^{*}(\mathbf{s}_{h},a_{L})-Q_{h}^{*}(\mathbf{s}_{h},a_{S})=$ $\beta\log\frac{\pi_{h}^{*}(a_{L}|\mathbf{s}_{h})}{\mu(a_{L})}-\beta\log\frac{\pi_{h}^{*}(a_{S}|\mathbf{s}_{h})}{\mu(a_{S})}$
$=$ $\beta\log\frac{\pi_{h}^{*}(a_{L}|\mathbf{s}_{h})}{\pi_{h}^{*}(a_{S}|\mathbf{s}_{h})}-\beta\log\frac{\mu(a_{L})}{\mu(a_{S})},$ (4)
@@
$\Pr(a_{S}\succ a_{L}|\mathbf{s}_{h})$ $=\frac{1}{1+\frac{\pi_{h}^{*}(a_{L}|\mathbf{s}_{h})}{\pi_{h}^{*}(a_{S}|\mathbf{s}_{h})}}=\pi_{h}^{*}(a_{S}|\mathbf{s}_{h}),$
where the latter equation is due to the fact that $\pi_{h}^{*}(a_{S}|\mathbf{s}_{h})+\pi_{h}^{*}(a_{L}|\mathbf{s}_{h})=1$. Therefore, given a sequence of token $\mathbf{s}_{h}$, once we have labeled the preference $\operatorname{\mathds{1}}[a_{S}\succ a_{L}|\mathbf{s}_{h}]$, $\pi_{h}(a_{S}|\mathbf{s}_{h})$, the routing agent $\pi$ can be learned by minimizing the cross-entropy loss
@@
where $\operatorname{\mathds{1}}[a_{L}\succ a_{S}|\mathbf{s}_{h}]$ indicates using large language model is preferred at state $\mathbf{s}_{h}$.
2.3 Acquiring Token-level Routing Preference
In this subsection, we describe our strategy to determine the preference label $\operatorname{\mathds{1}}[a_{L}\succ a_{S}|\mathbf{s}_{h}]$. For a state $\mathbf{s}_{h}$, we first generate the next token $y_{h+1}$ with the small language model and then complete the whole trajectory $\mathbf{s}_{H}$ until the generation ends with <EOS> using the routing policy $\pi$. Compared to equation 1, the reward collected on this trajectory $\mathbf{s}_{H}$ is an unbiased estimation of $Q_{h}^{\pi}(\mathbf{s}_{h},a_{S})$. Intuitively, if using the small language model in the current step $h$ can generate the correct response $\mathbf{s}_{H}$, then the small language model is preferred. Otherwise, we assign $a_{L}\succ a_{S}$ and assume that the large language model can generate the correct answer, as implemented in Line 9 in Algorithm 1.
However, generating and evaluating the final response $\mathbf{s}_{H}$ might be expensive or even infeasible when $H$ is large. In order to further accelerate the token-level routing preference label, we introduce a shortcut by leveraging the ground truth response $\mathbf{s}^{*}_{H}$ provided in the dataset. As Line 8 in Algorithm 1 suggests, if the next token $y_{h+1}^{S}$ generated by the small language model is exactly the same as the ground-truth token $y^{*}_{h+1}$, we assign $a_{S}\succ a_{L}$ since the behavior of the small language model is good enough to match the ground-truth model. In the case where the next token generated by the small language model does not match the ground truth, as carried out in Line 10, we check the next token generated by the large language model $y_{h+1}^{L}$ and assign $a_{L}\succ a_{S}$ if $y_{h+1}^{L}=y_{h+1}^{*}$. Otherwise we will go back to the aforementioned case to evaluate the completed generated trajectory as conducted in Line 10 in Algorithm 1. We would like to highlight that only when both models fail to generate the correct token $y^{*}_{h}$ based on ground truth context, the full response generation is required to compute the reward. This shortcut allows us to obtain routing preferences for most tokens without generating the full response. Empirically, we find that about $80\%\sim 90\%$ of the tokens can be correctly predicted by either the SLM or LLM, which makes the shortcut significantly reduce the inference costs of estimating the reward function.
@@ 2.3 Acquiring Token-level Routing Preference
8:    if $M_{S}(\mathbf{s}_{h})=y^{*}_{h+1}$ then Set $p_{h}=1$ /* Case 1. $a_{S}\succ a_{L}$ */
9:    if $M_{L}(\mathbf{s}_{h})=y^{*}_{h+1}$ then Set $p_{h}=0$ /* Case 2. $a_{L}\succ a_{S}$ */
@@
2.4 Proposed Algorithm
The inference algorithm CITER is presented in Algorithm 2. In detail, CITER uses a deterministic policy where it chooses the small language model when $\pi(a_{S}|\mathbf{s})\geq\pi(a_{L}|\mathbf{s})$ (i.e., $\pi(a_{S}|\mathbf{s})\geq\frac{1}{2}$) and vice versa. To further investigate the balance between efficiency and precision by collaborating with LLM and SLM, we introduce another layer of prior policy $(\rho(a_{S}),\rho(a_{L}))$, where $\rho(a_{S})+\rho(a_{L})=1$. Thus, the deterministic rule of selecting the SLM from the posterior policy distribution $\pi^{\prime}(a|\mathbf{s})\propto\pi(a|\mathbf{s})\rho(a)$ is that
Remark 2.1.
We maintain separate KV caches for SLM and LLM. When CITER switches between them, the previous KV cache is preserved, allowing it to be reused when switching back. This eliminates the need for redundant computations, improving efficiency.
@@ 3 Experiments
In this section, we evaluate the performance of CITER aiming to answer the following questions: (1) Compared with the previous works on speeding up the inference of LLM, how does our framework perform in terms of the inference costs and the response quality? (2) Does the components we proposed in our framework boost the performance of the router? (3) Does the iterative training process of the router improve the performance of our framework? (4) How does the performance of our framework change with the size of the LLM? (5) Can the router distinguish the critical and non-critical tokens correctly?
@@ 3.1 Experimental Setup
Table 1: The statistics of our evaluation datasets. The commonsense QA dataset and MMLU-Professional Psychology dataset are denoted as CS QA and MMLU-$P$, respectively.
@@
MMLU-$P$ Psychology Multi-choice 4 612 69
@@
Evaluation. To evaluate the performance of CITER and the baseline methods, we use the accuracy of responses to reflect response quality and define the inference cost as the average amount of data transformation per sample, mainly the KV cache that must be transferred from GPU HBM to the on-chip cache, since LLM generation is primarily memory-bound. Details on the data transformation calculations and an illustration of the memory-bound nature of LLM generation are provided in Appendix E. Additionally, for both query-level routing methods and token-level routing methods, a threshold $\tau$ is applied in each method to balance the trade-off between leveraging the LLM to improve the response quality or offloading to the SLM to reduce the overall inference costs. We then plot the accuracy curve versus the average amount of data transformation per sample to illustrate the acceleration performance of both CITER and the baselines. The optimal point is located in the top-left corner of the curve, corresponding to the highest accuracy with the lowest costs.
@@
Implementation Details We implement our framework using the Hugging Face Transformers library (Wolf et al., 2020). For the SLM and LLM, we utilize Qwen2-1.5b and Qwen2-72b, respectively. The router is implemented as a multilayer perceptron (MLP) network with three hidden layers, ReLU activation (Agarap, 2019), BatchNorm normalization (Ioffe & Szegedy, 2015), and a $0.1$ dropout rate. It is trained using the Adam optimizer (Kingma & Ba, 2017) with a learning rate of $1\times 10^{-7}$, betas of $(0.9,0.99)$, and no weight decay. Training is performed on a single NVIDIA H100 GPU with a batch size of $80$. The iterative training process runs for $2$ rounds. We use the hidden state corresponding to the last generated token from the SLM as the input to our router. This approach enables the router to utilize the rich representations extracted by the SLM, allowing routing decisions to be informed not only by the current token but also by the broader context accumulated thus far.